# Conditioned quantum-assisted deep generative surrogate for particle-calorimeter interactions

### v1

@@ III Results
Table 1: Fréchet Particle Distance (FPD) and Kernel Particle Distance (KPD) metrics, implemented in the JetNet [56] library and adapted for the CaloChallenge [19, 57].
@@
MCMC (100k) $(442.98\pm 2.44)\times 10^{-3}$ $(0.84\pm 0.10)\times 10^{-3}$
Training: We train our model for $220$ epochs. Each epoch typically has $625$ updating steps. The number of block Gibbs sampling steps was set to $3000$. During the first $100$ epochs we anneal the model parameters, such as those used in the Gumbel trick, from smooth to a sharp step, to mitigate the discontinuities in the gradient related to the use of discrete variables. After the annealing parameters have reached their final values, we train the model for the next $50$ epochs, after which we freeze the encoder and decoder parameters at epoch $150$, while the prior distribution parameters keep being updated. We use an Nvidia A100 GPU. Our experiment show that this last step was necessary for the RBM log-likelihood to saturate. In Fig. 6 we show the RBM log-likelihood vs epochs. The yellow star point to when the annealed parameters reached their final values, whereas the purple star point to when the encoder and decoder parameter were frozen, after which it is clear the log-likelihood saturates. To estimate the log-likelihood, after training, we used annealed importance sampling and reverse annealed importance sampling methods [58, 59] with a annealed step set to $1/30$. As the model updates its parameters, the encoded data used to estimate the log-likelihood will also be modified from epoch to epoch, hence, we stored the validation encoded data for each epoch in order to accurately estimate the RBM log-likelihood. During training we save an instance of the model every ten epochs.
@@
Validation: To validate our model, we use Fréchet physics distance (FPD) and the kernel physics distance (KPD) [21]. These are integral probability metrics for high energy physics, specifically designed to be sensitive to modelling of shower shape variables. In Fig. 6 we show these metrics vs epochs. Henceforth, the results are generated using the $220$-epoch model, corresponding to the best FPD and KPD metrics, together with a saturated RBM log-likelihood. We further validate our model’s reconstruction, classical sampling and QA sampling with Geant4 data. To sample using the Advantage_system6.4 QA [25], we first estimate the effective $\beta$ by iteratively using Eq. (23) until the absolute difference between the QA’s energy and that of the RBM is smaller than the reduced standard error, viz.,
@@
In the case of conditioned sampling using the QA, we estimate the temperature under flux biases. We observed that using flux biases increases the QA’s effective temperature. Additionally, programming the QA increases temperature fluctuations. In Fig. 8 (a-c) we present the RBM energy histogram obtained from encoded Geant4 showers, classically generated samples and QA-generated showers. Specifically, in Fig. 8 a) we estimate the inverse temperature for each incident energy condition, as outlined previously, before generating the QA sample. Fig. 8 d) shows the estimated inverse temperature per incident energy condition. Notably, we observe rather large but rare thermal fluctuations of the order of $10\%$. In Fig. 8 e) we show the same estimated inverse temperature shown in Fig. 8 d) versus the incident energy condition, suggesting that these large thermal fluctuations are independent of the incident energy condition. To further investigate, we repeated the process of generating conditioned samples using the QA, but estimated the QA inverse temperature only once at the beginning. This estimate is depicted as a dashed black line Fig. 8 d). In Fig. 8 b) we show the RBM energy, where a small shift in the energy of the QA samples is noticeable. We repeated the previous setting and introduced a $2.5$s pause between QA API calls. In Fig. 8 c) we show the RBM energy obtained from the samples, from which it is clear that the energy shift has disappeared. Additionally, in Fig. 8 d) we show the estimated inverse temperature in the absence of flux biases in dashed red. We conclude that there are two main contributions to the thermal fluctuations: the programming of the QA and the flux biases. The former can be substantially mitigated by either pausing the sampling process between samples or estimating the inverse temperature per sample. The latter can be accounted for by estimating the inverse temperature in the presence of flux biases. We discuss this further in the Discussion section.
@@
After estimating the effective $\beta$ in the QA, we use the validation dataset composed of $10,000$ data points to benchmark the model. In Fig. 9 (a) we show the shower energy histograms, while in Fig. 9 (b) we show the sparsity histogram. The sparsity index is a measure of the sparsity of the shower and we define it as the ratio between zero-value voxels in the shower and total number of voxels. On each of the histograms we include the Geant4 data and the model’s reconstruction. Furthermore, we test the generative model by generating samples from the RBM and feeding these samples together with the incidence energy from the validation dataset to the decoder and generating shower samples. Similarly, we test the quantum-assisted generative model by generating samples from the QA instead of the classical RBM. Both, the classically generated samples and the quantum-assisted generated ones are included in Figs. 9 (a) and (b) and are labelled Sample and Sample w/ QPU, respectively. In Fig. 9 (c) we show the RBM energy distribution corresponding to the encoded validation dataset and, both, the classically and QA sampled data.
@@
To further validate our model, we generate a synthetic dataset composed by $100,000$ events which the same incidence energy distribution as the training dataset. We compare our synthetic dataset with Dataset 2’s test set [3] by computing the mean energy in the $r,\theta$ and $z$ directions of the cylinder (see Fig. 1). We show these results in Figs. 9 (d-f).
@@ IV Discussion
In the previous section we outlined the process of training our conditioned quantum-assisted Calo4pQVAE and described a novel a faster method to estimate the QA’s effective inverse temperature before sampling using the QPU. We evaluated the performance of our model by means of the KPD and FPD metrics, achieving results on the same order and one order of magnitude higher than CaloDiffusion, respectively, as shown in Table 1 [19]. Our results demonstrate that i) our framework is able of reconstructing the showers and preserving sparsity, and ii) the prior is effectively learning the structure of the encoded data.
@@
We previously mentioned that we set the number of block Gibbs sampling steps to 3000. We observed that with this number of steps, the RBM log-likelihood versus epochs increased monotonically. In this case, generating $1024$ samples classically takes approximately $1$ s with $O(1000)$ MCMC steps. Conversely, when sampling with QA, there are three characteristic timescales to consider: the programming time (approximately $10$ ms), the annealing time (approximately $20\mu s$) and the readout time (approximately $100\mu s$). This results in a total generation time of approximately $0.1$ s for O(1000) samples, i.e., one order of magnitude faster than the classical method, assuming the QPU programming step is performed only once. These estimates per sample are shown in Table 2. However, in our current framework, as mentioned earlier, the flux biases conditioning is done during the programming step, effectively increasing the overall time to approximately $10$s.
@@ Appendix E Quadripartite RBM numerical verification
We considered different number of nodes ranging from … for 4-partite RBMs. This setting allows for the explicit enumeration of all feasible states, facilitating the precise computation of the partition function. To rigorously assess the accuracy of our approach, we conducted a comparative analysis of the density of energy states. This entailed a direct comparison between the utilization of all feasible states and the implementation of the 4-partite Gibbs sampling method, as elaborated upon in Section II.3. In Figure E1, we present a detailed visual comparison of the density of states obtained through both methodologies, across various iterations of the Gibbs sampling process. This comparison shows the convergence and consistency of these two approaches.

### v2

@@ III Results
Table 1: Fréchet Particle Distance (FPD) and Kernel Particle Distance (KPD) metrics, implemented in the JetNet [57] library and adapted for the CaloChallenge [3].
@@
MCMC (100k) $(390.35\pm 1.85)\times 10^{-3}$ $(0.46\pm 0.05)\times 10^{-3}$
Training: We train our model for $250$ epochs. Each epoch typically has $625$ updating steps. The number of block Gibbs sampling steps was set to $3000$. During the first $50$ epochs we anneal the model parameters, such as those used in the Gumbel trick, from smooth to a sharp step, to mitigate the discontinuities in the gradient related to the use of discrete variables. After the annealing parameters have reached their final values, we train the model for the next $150$ epochs, after which we freeze the encoder and decoder parameters at epoch $200$, while the prior distribution parameters keep being updated. We use an Nvidia A100 GPU. Our experiments show that this last step was necessary for the RBM log-likelihood to saturate. In Fig. 6 we show the RBM log-likelihood vs epochs. The yellow star points to when the annealed parameters reached their final values, whereas the purple star points to when the encoder and decoder parameter were frozen, after which it is clear the log-likelihood saturates. To estimate the log-likelihood, after training, we used annealed importance sampling and reverse annealed importance sampling methods [58, 59] with an annealed step set to $1/30$. As the model updates its parameters, the encoded data used to estimate the log-likelihood will also be modified from epoch to epoch, hence, we stored the validation encoded data for each epoch in order to accurately estimate the RBM log-likelihood. During training we save an instance of the model every ten epochs.
@@
Validation: To validate our model, we use Fréchet physics distance (FPD) and the kernel physics distance (KPD) [22]. These are integral probability metrics for high energy physics, specifically designed to be sensitive to modelling of shower shape variables. In Fig. 6 we show these metrics vs epochs. Henceforth, the results are generated using the $240$-epoch model, corresponding to the best KPD metric with good FPD metric in the saturated RBM log-likelihood regime. We further validate our model’s reconstruction, classical sampling and QA sampling with Geant4 data. To sample using the Advantage_system6.4 QA [26], we first estimate the effective $\beta$ by iteratively using Eq. (23) until the absolute difference between the QA’s energy and that of the RBM is smaller than the reduced standard error, viz.,
@@
In the case of conditioned sampling using the QA, we estimate the temperature under flux biases. We observed that using flux biases increases the QA’s effective temperature. Additionally, programming the QA increases temperature fluctuations. In Fig. 8 (a-c) we present the RBM energy histogram obtained from encoded Geant4 showers, classically generated samples and QA-generated showers. Specifically, in Fig. 8 a) we estimate the inverse temperature for each incident energy condition, as outlined previously, before generating the QA sample. Fig. 8 d) shows the estimated inverse temperature per incident energy condition. Notably, we observe rather large but rare thermal fluctuations of the order of $10\%$. In Fig. 8 e) we show the same estimated inverse temperature shown in Fig. 8 d) versus the incident energy condition, suggesting that these large thermal fluctuations are independent of the incident energy condition. To further investigate these effects, we repeated the process of generating conditioned samples using the QA, but estimated the QA inverse temperature only once at the beginning of the sampling process. This estimate is depicted as a dashed black line Fig. 8 d). In Fig. 8 b) we show the RBM energy, where a small shift in the energy of the QA samples is noticeable. Using the same inverse temperature obtained in the previous setting, we generated new samples and introduced a $2.5$s pause between QA API calls, that is per sample. In Fig. 8 c) we show the RBM energy obtained from these new samples, from which it is clear that the energy shift has disappeared. Additionally, in Fig. 8 d) we show the estimated inverse temperature in the absence of flux biases in dashed red. We conclude that there are two main contributions to the thermal fluctuations: the programming of the QA and the flux biases. The former can be substantially mitigated by either pausing the sampling process between samples or estimating the inverse temperature per sample. The latter can be accounted for by estimating the inverse temperature in the presence of flux biases. We discuss this further in the Discussion section.
@@
After estimating the effective $\beta$ in the QA, we use the validation dataset composed of $10,000$ data points to benchmark the model. In Fig. 9 (a) we show the shower energy histograms, while in Fig. 9 (b) we show the sparsity histogram. The sparsity index is a measure of the sparsity of the shower and we define it as the ratio between zero-value voxels in the shower and total number of voxels. Each of the histograms include the Geant4 data and the model’s reconstruction. Furthermore, we test the generative model by generating samples from the RBM and feeding these samples together with the incidence energy from the validation dataset to the decoder and generating shower samples. Similarly, we test the quantum-assisted generative model by generating samples from the QA instead of the classical RBM. Both, the classically generated samples and the quantum-assisted generated ones are included in Figs. 9 (a) and (b) and are labelled Sample and Sample w/ QPU, respectively. In Fig. 9 (c) we show the RBM energy distribution corresponding to the encoded validation dataset and, both, the classically and QA sampled data.
@@
To further validate our model, we generate a synthetic dataset composed by $100,000$ events with the same incidence energy distribution as the training dataset. We compare our synthetic dataset with Dataset 2’s test set [3] by computing the mean energy in the $r,\theta$ and $z$ directions of the cylinder (see Fig. 1). We show these results in Figs. 9 (d-f).
@@ IV Discussion
In the previous section we outlined the process of training our conditioned quantum-assisted Calo4pQVAE and described a novel and faster method to estimate the QA’s effective inverse temperature before sampling using the QPU. We evaluated the performance of our model by means of the KPD and FPD metrics, achieving results on the same order and one order of magnitude higher than CaloDiffusion, respectively, as shown in Table 1 [19]. Furthermore, when compared with the models in the CaloChallenge under this metric, our framework performs better than more than half of the 18 models considered [21]. Our results demonstrate that i) our framework is able of reconstructing the showers and preserving sparsity, and ii) the prior is effectively learning the structure of the encoded data.
@@
We previously mentioned that we set the number of block Gibbs sampling steps to 3000. We observed that with this number of steps, the RBM log-likelihood versus epochs increased monotonically. In this case, generating $1024$ samples classically takes approximately $1$ s with $O(1000)$ MCMC steps. Conversely, when sampling with QA, there are three characteristic timescales to consider: the programming time (approximately $10$ ms), the annealing time (approximately $20\mu s$) and the readout time (approximately $100\mu s$). This results in a total generation time of approximately $0.1$ s for O(1000) samples, i.e., one order of magnitude faster than the classical method, assuming the QPU programming step is performed only once. These estimates per sample are shown in Table 2, which are highly competitive [21]. However, in our current framework, as mentioned earlier, the flux biases conditioning is done during the programming step, effectively increasing the overall time to approximately $10$s.
@@ Appendix E Quadripartite RBM numerical verification
We considered 4-partite RBM with six nodes per partition. This setting allows for the explicit enumeration of all feasible states, facilitating the precise computation of the partition function. To rigorously assess the accuracy of our approach, we conducted a comparative analysis of the density of energy states. This entailed a direct comparison between the utilization of all feasible states and the implementation of the 4-partite Gibbs sampling method, as elaborated upon in Section II.3. In Figure E1, we present a detailed visual comparison of the density of states obtained through both methodologies, across various iterations of the Gibbs sampling process. This comparison shows the convergence and consistency of these two approaches.